Ignition wires heat sample Thermometer Stirrer A bomb calorimeter, or constant volume calorimeter, is a device often used to determine the heat of combustion of fuels and the energy content of foods. Since the "bomb" itself can absorb energy, a separate experiment is needed to determine the heat capacity of the calorimeter This is known as calibrating the calorimeter In the laboratory a student burns a 0.319-g sample of phenanthrene (C14H10 in a bomb calorimeter containing 1070. g of water. The temperature increases from 25.60 °C to 28.00 °C. The heat capacity of water is 4.184J gl°c-1 The molar heat of combustion is-7054 kJ per mole of phenanthrene Water C14Hios) +33/2 02(g) 14 CO2(8)+5 H20)Energy Calculate the heat capacity of the calorimeter Insulated Sample Bring outside chamber heat capacity of calorimeter = dish sample bomb Combustion (bomb) calorimeter
Thermometer Ignition wires heat sample Stirrer A bomb calorimeter, or a constant volume calorimeter, is a device often used to determine the heat of combustion of fuels and the energy content of foods. In an experiment, a 0.5274 g sample of diphenylacetylene (C14H is burned completely in a bomb calorimeter. The calorimeter is surrounded by 1.348x103 g of water. During the combustion the temperature increases from 21.86 to 25.00 C. The heat capacity of water is 4.184 Jgloc1 The heat capacity of the calorimeter was determined in a p Water revious experiment to be 963.1 J/°C. Assuming that no energy is lost to the surroundings, calculate the molar heat of combustion of diphenylacetylene based on these data. C14H10s)+(33/2) 02(8)>5 H20()+14 CO2(g)+Energy Molar Heat of Combustion Insulated Sample Burning outside chamber kJ/mol dish sample bomb Combustion (bomb) calorimeter
Answer #1

Question 1:

Number of moles of Phenathrene = Mass/Molar mass = 0.319/178.2 = 0.00179 moles

Heat evolved due to combustion = (Molar heat of combustion) * (Number of moles) = 0.00179 mol * 7054 kJ/mol =12627 J

Heat used by water = mass of water * specific heat capacity of water * Change in temperature

=> 1070 * 4.184 * (28.00 - 25.60)

=> 10744 J

Heat gained by calorimeter = 12627 - 10744 = 1883 J

Heat capacity of calorimeter = (Heat gained by calorimeter)/(Change in temperature) = 1883/(28.00-25.60) = 784.5 J/C

Question 2:

Heat gained by water = mass of water * specific heat capacity of water * Change in temperature

=> 1348* 4.184 * (25.00 - 21.86)

=> 17709.7 J

Heat gained by calorimeter = (963.1 J/C) * (25.00-21.86)C = 3024.1 J

Heat released by combustion = 17709.7 + 3024.1 = 20733.8 J

Number of moles of Diphenylacetylene = Mass/Molar mass = 0.5274/178.2 = 0.002959 moles

Molar Heat of combustion = -20733.8/0.002959 = -7007 kJ/mol

Note - Post any doubts/queries in comments section.